OpenStudy (anonymous):

Find (a) the equation of a line that is parallel to the given line and includes the given point, and (b) the equation of a line that is perpendicular to the given line through the given point. Write both answers in slope-intercept form. (c) Graph both of these lines on the same axes. Please show all of your work. y = 2x + 6, (-1, -2)

OpenStudy (anonymous):

y = 2x + 6, (-1, -2) here the slope m = 2

OpenStudy (anonymous):

next use the formula y-y1=m(x-x1)

OpenStudy (anonymous):

for parallel line use m=m but for perpendicukar line of it use m=-1/m

OpenStudy (anonymous):

for this point (-1, -2) the parallel line is y-y1=m(x-x1) y-(-2)=m(x-(-1)) y+2=2(x+1)

OpenStudy (anonymous):

for the perpendicular line of y=2x+6 we use m=-1/2 we have y-y1=m(x-x1) y-(-2)=m(x-(-1)) y+2=(-1/2)(x+1)

OpenStudy (anonymous):

now mult and add y+2=2(x+1) y=-2+2x+2 =2x ans....this is the parallel line of y=2x+6

OpenStudy (anonymous):

for the perpendicular line of y=2x+6 at point(-1,-2) we use m=-1/2 we have y-y1=m(x-x1) y-(-2)=m(x-(-1)) here use m=-1/2 y+2=(-1/2)(x+1)

OpenStudy (anonymous):

now mult and add y+2=-1/2 (x+1) y=-2-1/2x+2 =-1/2x ans....this is the perpendicular line of y=2x+6 at point(-1.-2)

OpenStudy (anonymous):

Thank you so much Mark O.